Two charges, -8 C and -4C are positioned 10 meters from each other. What does the -8C feel as a result of the -4C? 


1
Expert's answer
2022-02-21T14:56:10-0500

We know that

Repulsion force

"F=\frac{kq_1q_2}{r^2}"

"F=\frac{9\times10^{9}\times(-8)(-4)}{10^2}=2.88\times10^{9}N"

Feel repulsion force